# Two-stage Estimation for Causal Inference Involving a Semi-continuous Exposure

### v1

@@ 2.1 Notation and Assumptions
To identify the causal effects we extend the standard causal inference assumptions (Rubin, 1980, 1990; Rosenbaum and Rubin, 1983; Cole and Hernán, 2008) to the setting of a semi-continuous exposure; Web Appendix A provides a detailed discussion.
@@ 2.3 A New Marginal Structural Model (MSM)
where $Y(A,D)$ is the potential outcome, $(A,D)\perp Q$, $Q$ has a mean of zero, and $\boldsymbol{\psi}=(\psi_{0},\psi_{11},\psi_{12})^{\prime}$. Based on Assumption 3, the potential outcomes $Y(A,D)$ are independent of observed exposures $(A,D)$ given covariates $\mathbf{X}$ (Rosenbaum and Rubin, 1983) which implies $Q$ is conditionally independent of $A$ and $D$. Under Assumptions 1-4, $\psi_{11}$ and $\psi_{12}$ represent the causal effects of the exposure at the mean dose and the causal dose-response effect; see Web Appendix B for more derivation details.
@@ 2.4 Regression Adjustment with Two Propensity Scores
where $\boldsymbol{\eta}=(\eta_{0},\boldsymbol{\eta}_{1}^{\prime},\boldsymbol{\eta}_{2}^{\prime})^{\prime}$ with $\boldsymbol{\eta}_{1}=(\eta_{11},\eta_{12})^{\prime}$ and $\boldsymbol{\eta}_{2}=(\eta_{21},\eta_{22})^{\prime}$. Under Assumptions 14 and correct specification of the exposure models, conditioning on $\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})$ balances covariates, and $\eta_{11}$ and $\eta_{12}$ identify the causal effect of exposure at the reference dose and the causal dose–response effect, respectively. Detailed derivations are given in Web Appendix C.
@@ 3.3.2 Large Sample Theory
Proofs of Theorems 13 and explicit expressions for the sandwich covariance estimator and Wald statistics are given in Web Appendix D. We use the resulting sandwich covariance matrix for $\hat{\boldsymbol{\Omega}}$ to compute standard errors and construct Wald tests and confidence intervals.
@@ 4.1 Misspecification of the Stage I Propensity Score
where $\rho_{1}=\text{corr}\{E(D|\mathbf{X},A=1),\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})\}$, $\boldsymbol{\zeta}_{1}(\mathbf{X}|A=1)=\{\zeta_{11}(\mathbf{X}|A=1),\ldots,\zeta_{1k}(\mathbf{X}|A=1)\}^{\prime}$, and $\boldsymbol{\beta}_{1}(\mathbf{X}|A=1)=\{\beta_{11}(\mathbf{X}|A=1),\ldots,\beta_{1k}(\mathbf{X}|A=1)\}^{\prime}$, with $\zeta_{1j}(\mathbf{X}|A=1)=\text{cov}\{E(D|\mathbf{X},A=1),\mathbf{X}_{j}|A=1\}$ and $\beta_{1j}(\mathbf{X}|A=1)=\text{cov}\{\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1}),\mathbf{X}_{j}|A=1\}$ for $j=1,\ldots,k$. Further details are provided in Web Appendix E.1.
@@ 4.2 Biased Estimator in Stage II
Further details are provided in Web Appendix E.2.
@@ 5.2 Methods
To examine robustness, we consider five misspecification scenarios: (i) misspecified Stage I PS $S_{1}$; (ii) misspecified Stage II PS $S_{2}$; (iii) misspecified imputation model $m_{a}(\mathbf{X};\boldsymbol{\theta})$; (iv) misspecification of $S_{1}$ and $m_{a}(\mathbf{X};\boldsymbol{\theta})$; and (v) misspecification of $S_{2}$ and $m_{a}(\mathbf{X};\boldsymbol{\theta})$. In all cases we focus on the setting with minimal propensity scores $\mathbf{S}=\{S_{1}(\mathbf{Z}_{1}),S_{2}(\mathbf{Z}_{2})\}^{\prime}$. Full details of how each model is misspecified are given in Web Appendix F.1.
@@ 5.3 Finite Sample Performance
Web Tables 1–3 and 4–6 in Web Appendix F.2 present results for $P(A=1)=0.25$ and $0.75$ respectively, and show patterns consistent with those for $P(A=1)=0.5$.
@@ 6.1 Description of the data
We adjust for baseline maternal sociodemographic characteristics, smoking and other substance use during pregnancy, reproductive history, gestational age at screening, and measures of the home environment and maternal cognition (listed in Web Table 7 in Web Appendix G.1). To address missingness, we generate 20 multiply imputed datasets using an imputation model including all variables; further details are given in Web Appendix G.2. Point estimates and standard errors are combined using Rubin’s rules under a missing-at-random assumption (Rubin, 1976; Little and Rubin, 2019).
@@ 6.2 Application and findings
We estimate the causal effects of prenatal drinking status and continuous dose using five approaches: (i) conventional covariate regression adjustment; (ii) two–propensity-score regression adjustment; (iii) a two-stage analysis with PS regression adjustment in Stage II; (iv) a two-stage analysis with IPW in Stage II; and (v) a two-stage analysis with AIPW in Stage II. Implementation details are provided in Web Appendix G.3.
Web Table 7 summarises the fitted models for the continuous and binary exposure components, $E(D|A=1,\mathbf{X})$ and $P(A=1|\mathbf{X})$. Both models include all candidate covariates, following the recommendation of Brookhart et al. (2006) that including variables related to the outcome can improve precision of causal effect estimates.

### v2

@@ 2.1 Notation and Assumptions
To identify the causal effects we extend the standard causal inference assumptions (Rubin, 1980, 1990; Rosenbaum and Rubin, 1983; Cole and Hernán, 2008) to the setting of a semi-continuous exposure; Appendix A provides a detailed discussion.
@@ 2.3 A New Marginal Structural Model (MSM)
where $Y(A,D)$ is the potential outcome, $(A,D)\perp Q$, $Q$ has a mean of zero, and $\boldsymbol{\psi}=(\psi_{0},\psi_{11},\psi_{12})^{\prime}$. Based on Assumption 3, the potential outcomes $Y(A,D)$ are independent of observed exposures $(A,D)$ given covariates $\mathbf{X}$ (Rosenbaum and Rubin, 1983) which implies $Q$ is conditionally independent of $A$ and $D$. Under Assumptions 1-4, $\psi_{11}$ and $\psi_{12}$ represent the causal effects of the exposure at the mean dose and the causal dose-response effect; see Appendix B for more derivation details.
@@ 2.4 Regression Adjustment with Two Propensity Scores
where $\boldsymbol{\eta}=(\eta_{0},\boldsymbol{\eta}_{1}^{\prime},\boldsymbol{\eta}_{2}^{\prime})^{\prime}$ with $\boldsymbol{\eta}_{1}=(\eta_{11},\eta_{12})^{\prime}$ and $\boldsymbol{\eta}_{2}=(\eta_{21},\eta_{22})^{\prime}$. Under Assumptions 14 and correct specification of the exposure models, conditioning on $\mathbf{S}(\mathbf{X};\boldsymbol{\alpha})$ balances covariates, and $\eta_{11}$ and $\eta_{12}$ identify the causal effect of exposure at the reference dose and the causal dose–response effect, respectively. Detailed derivations are given in Appendix C.
@@ 3.3.2 Large Sample Theory
Proofs of Theorems 13 and explicit expressions for the sandwich covariance estimator and Wald statistics are given in Appendix D. We use the resulting sandwich covariance matrix for $\hat{\boldsymbol{\Omega}}$ to compute standard errors and construct Wald tests and confidence intervals.
@@ 4.1 Misspecification of the Stage I Propensity Score
where $\rho_{1}=\text{corr}\{E(D|\mathbf{X},A=1),\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1})\}$, $\boldsymbol{\zeta}_{1}(\mathbf{X}|A=1)=\{\zeta_{11}(\mathbf{X}|A=1),\ldots,\zeta_{1k}(\mathbf{X}|A=1)\}^{\prime}$, and $\boldsymbol{\beta}_{1}(\mathbf{X}|A=1)=\{\beta_{11}(\mathbf{X}|A=1),\ldots,\beta_{1k}(\mathbf{X}|A=1)\}^{\prime}$, with $\zeta_{1j}(\mathbf{X}|A=1)=\text{cov}\{E(D|\mathbf{X},A=1),\mathbf{X}_{j}|A=1\}$ and $\beta_{1j}(\mathbf{X}|A=1)=\text{cov}\{\tilde{S}_{1}(\mathbf{X};\tilde{\boldsymbol{\alpha}}_{1}),\mathbf{X}_{j}|A=1\}$ for $j=1,\ldots,k$. Further details are provided in Appendix E.1.
@@ 4.2 Biased Estimator in Stage II
Further details are provided in Appendix E.2.
@@ 5.2 Methods
To examine robustness, we consider five misspecification scenarios: (i) misspecified Stage I PS $S_{1}$; (ii) misspecified Stage II PS $S_{2}$; (iii) misspecified imputation model $m_{a}(\mathbf{X};\boldsymbol{\theta})$; (iv) misspecification of $S_{1}$ and $m_{a}(\mathbf{X};\boldsymbol{\theta})$; and (v) misspecification of $S_{2}$ and $m_{a}(\mathbf{X};\boldsymbol{\theta})$. In all cases we focus on the setting with minimal propensity scores $\mathbf{S}=\{S_{1}(\mathbf{Z}_{1}),S_{2}(\mathbf{Z}_{2})\}^{\prime}$. Full details of how each model is misspecified are given in Appendix F.1.
@@ 5.3 Finite Sample Performance
Tables 1–3 and 4–6 in Appendix F.2 present results for $P(A=1)=0.25$ and $0.75$ respectively, and show patterns consistent with those for $P(A=1)=0.5$.
@@ 6.1 Description of the data
We adjust for baseline maternal sociodemographic characteristics, smoking and other substance use during pregnancy, reproductive history, gestational age at screening, and measures of the home environment and maternal cognition (listed in Table 7 in Appendix G.1). To address missingness, we generate 20 multiply imputed datasets using an imputation model including all variables; further details are given in Appendix G.2. Point estimates and standard errors are combined using Rubin’s rules under a missing-at-random assumption (Rubin, 1976; Little and Rubin, 2019).
@@ 6.2 Application and findings
We estimate the causal effects of prenatal drinking status and continuous dose using five approaches: (i) conventional covariate regression adjustment; (ii) two–propensity-score regression adjustment; (iii) a two-stage analysis with PS regression adjustment in Stage II; (iv) a two-stage analysis with IPW in Stage II; and (v) a two-stage analysis with AIPW in Stage II. Implementation details are provided in Appendix G.3.
Table 7 in Appendix G.1 summarises the fitted models for the continuous and binary exposure components, $E(D|A=1,\mathbf{X})$ and $P(A=1|\mathbf{X})$. Both models include all candidate covariates, following the recommendation of Brookhart et al. (2006) that including variables related to the outcome can improve precision of causal effect estimates.